# Knowledge Distillation of a Protein Language Model Yields a Foundational Implicit Solvent Model

### v1

@@ 3 Introduction
These deficiencies stem from two overarching limitations. First, ISMs rely on approximate analytical expressions to compute the solvation free energy, $E_{\mathrm{solv}}$; such approximations cannot fully capture the complex dependence of $E_{\mathrm{solv}}$ on molecular composition, geometry, and conformational state. Second, the parameters of ISMs are rarely optimized in a systematic or data-driven manner to reproduce results from experiment or explicit solvent simulations across diverse protein families. The rise of machine learning (ML) offers a path to overcome these challenges. By replacing the approximate analytical formulas with flexible neural network potentials, ML-based force fields could lead to ISMs with much-improved accuracy. Many promising studies have demonstrated the potential of using ML methodologies for force field development.airas_transferable_2023, katzberger_implicit_2023, katzberger_general_2024, katzberger_rapid_2025, katzberger_transferring_2025, chen_machine_2021, husic_coarse_2020, kozinsky_scaling_2023, wang_improving_2023, brunken_machine_2024, yao_machine_2023, wang_machine_2019, wang_multi-body_2021, fu_forces_2023, durumeric_learning_2024, arts_two_2023, kohler_flow-matching_2023, loose_coarse-graining_2023, durumeric_adversarial-residual-coarse-graining_2019, durumeric_machine_2023, ding_contrastive_2022, ding_optimizing_2024, chmiela_towards_2018, chmiela_accurate_2020, duschatko_thermodynamically_2024, ggreener_differentiable_2024, yang_slicing_2023, corso_graph_2024, batzner_advancing_2023, eastman_openmm_2024, zheng_predicting_2024, sahrmann_utilizing_2023, bonneau_peering_2025, anstine_machine_2023, cheng_developing_2024, faller_density-based_2024, takaba_machine-learned_2024, wang_espalomacharge_2024, wang_end--end_2022, galvelis_nnpmm_2023, sabanes_zariquiey_enhancing_2024, barnett_neural_2024, yang_construction_2022, charron_navigating_2025, majewski_machine_2023, liao_development_2025, riveros_neat-dna_2025, smith_ani-1_2017, devereux_extending_2020, gao_torchani_2020, unke_biomolecular_2024, zhang_pretraining_2024, klein_operator_2025, plainer_consistent_2025, murtada_md-llm-1_2025
@@ 6 Methods
Because Schake operates on backbone atoms rather than residues, each residue-level SS8 probability or label was duplicated for the corresponding $\alpha, C$, and N atoms. As a result, both $\bm{P}^{\mathrm{ESM3}}_{k}$ and $\bm{Y}^{\mathrm{DSSP}}_{k}$ were expanded to dimension $(3n_{\mathrm{res}}\times 8)$ for ESM3 inference and GNN training.

### v2

@@ 3 Introduction
These deficiencies stem from two overarching limitations. First, ISMs rely on approximate analytical expressions to compute the solvation free energy, $E_{\mathrm{solv}}$; such approximations cannot fully capture the complex dependence of $E_{\mathrm{solv}}$ on molecular composition, geometry, and conformational state. Second, the parameters of ISMs are rarely optimized in a systematic or data-driven manner to reproduce results from experiment or explicit solvent simulations across diverse protein families. The rise of machine learning (ML) offers a path to overcome these challenges. By replacing the approximate analytical formulas with flexible neural network potentials, ML-based force fields could lead to ISMs with much-improved accuracy. Many promising studies have demonstrated the potential of using ML methodologies for force field development.airas_transferable_2023, katzberger_implicit_2023, katzberger_general_2024, katzberger_rapid_2025, katzberger_transferring_2025, chen_machine_2021, husic_coarse_2020, kozinsky_scaling_2023, wang_improving_2023, brunken_machine_2024, yao_machine_2023, wang_machine_2019, wang_multi-body_2021, fu_forces_2023, durumeric_learning_2024, arts_two_2023, kohler_flow-matching_2023, loose_coarse-graining_2023, durumeric_adversarial-residual-coarse-graining_2019, durumeric_machine_2023, ding_contrastive_2022, ding_optimizing_2024, chmiela_towards_2018, chmiela_accurate_2020, duschatko_thermodynamically_2024, greener_differentiable_2024, yang_slicing_2023, corso_graph_2024, batzner_advancing_2023, eastman_openmm_2024, zheng_predicting_2024, sahrmann_utilizing_2023, bonneau_peering_2025, anstine_machine_2023, cheng_developing_2024, faller_density-based_2024, takaba_machine-learned_2024, wang_espalomacharge_2024, wang_end--end_2022, galvelis_nnpmm_2023, sabanes_zariquiey_enhancing_2024, barnett_neural_2024, yang_construction_2022, charron_navigating_2025, majewski_machine_2023, liao_development_2025, riveros_neat-dna_2025, smith_ani-1_2017, devereux_extending_2020, gao_torchani_2020, unke_biomolecular_2024, zhang_pretraining_2024, klein_operator_2025, plainer_consistent_2025, murtada_md-llm-1_2025
@@ 6 Methods
Because Schake operates on backbone atoms rather than residues, each residue-level SS8 probability or label was triplicated for the corresponding $\alpha, C$, and N atoms. As a result, both $\bm{P}^{\mathrm{ESM3}}_{k}$ and $\bm{Y}^{\mathrm{DSSP}}_{k}$ were expanded to dimension $(3n_{\mathrm{res}}\times 8)$ for ESM3 inference and GNN training.